OpenStudy (anonymous):

A generator with an adjustable frequency of oscillation is wired in series to an inductor of L=2.50 mH and a capacitor of C=3.00 uF. At what frequency does the generator pro-duce the largest possible current amplitude in the circuit? (Answer: 1.84 kHz)

OpenStudy (anonymous):

At resonance frequency, the generator will produce the largest possible current amplitude in the circuit solution : wL = wC w = 1 divided by sqrt(LC) frequency (v) = 1 divided by 2*pi* sqrt(LC) on putting the values i m getting an answer 1.79 kHz

OpenStudy (anonymous):

base on the question, we know L=2.50 mH = 2.5 x 10^-3 H C=3.00 uF = 3 x 10 ^-6 F XL = XC w*L = 1/(wC) w^2*L*C = 1 w^2 = 1/LC w = sqrt(1/LC) w = sqrt(1/(2.5 x 10^-3 x 3 x 10^-6)) w = sqrt(1/(7.5x10^-9)) w = sqrt(1.3x10^8) w = 11547 w = 2*pi*f f = w/(2*pi) f =11547/(2*3.14) f = 11547/6.28 f = 1838 Hz f = 1.84 KHz @Idealist right???

OpenStudy (anonymous):

it was my horrible mistake the formula was wL = 1 divided by wC i had applied the right formula but out of curiosity i wrote wL=wC

OpenStudy (anonymous):

no problem @niksva ur right too., :)